OpenStudy (anonymous):

Using a number line, determine the midpoint of segment KL; K is at 9 and L is at 15. 1) 12 2) 13 3) 11

OpenStudy (anonymous):

Here, you solve it for me :D

OpenStudy (anonymous):

Can you list me the numbers between 9 and 15 including those numbers?

OpenStudy (anonymous):

~~~~~~~~~~~~~ ----|------|------|------- 9 12 15

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

Well, try what I asked. :P

OpenStudy (anonymous):

Write every number from 9 to 15 including 9 and 15.

OpenStudy (anonymous):

the formula is (a+b)/2 here a=9 b=15 so (a+b)/2=(9+15)/2=12

OpenStudy (anonymous):

9 10 11 12 13 14 15. im confused at wat to do after that